22. Write the equation of the parabola given the following information. Assume that the vertex is at the origin.
LR length of 12 and opens down
23. Write the equation of the parabola given the following information. Assume that the vertex is at the origin.
The distance from the focus to the directrix is 100

Answer by lwsshak3(11628)

22. Write the equation of the parabola given the following information. Assume that the vertex is at the origin.
LR length of 12 and opens down
..
Basic form of equation for parabola that opens down with vertex at origin:
x^2=-4py
LR=latus rectum=focal width=4p=12
p=3
Equation: x^2=-12y
..
23. Write the equation of the parabola given the following information. Assume that the vertex is at the origin.
The distance from the focus to the directrix is 100
p=50 (distance from vertex to focus or directrix on the axis of symmetry
4p=200
Assume parabola opens up:
Equation:x^2=200y